# DOEPatch: Dynamically Optimized Ensemble Model for Adversarial Patches Generation

### v1

@@ IV-E Ablation Studies
To validate the effectiveness of the proposed dynamically optimized ensemble model, we conducted the following two experiments compared to the Average Ensemble model. We designed ensemble models utilizing the one-stage algorithms YOLOv2 and YOLOv3 and obtained respective adversarial patches, DOEPatch(YY) and AEPatch(YY). Additionally, we conducted experiments targeting Faster R-CNN and YOLOv3, which yielded novel adversarial patches, AEPatch(YF) and DOEPatch(YF). The weight parameters, denoted by ${{\text{w}}_{0}}$ and ${{\text{w}}_{1}}$, represent the weights before the detection model, which are constantly optimized during training in the dynamically optimized ensemble model. We manually adjusted the weight parameters of the average ensemble model and conducted five sets of experiments on AEPatch(YF), marked with numerical identifiers, as shown in Tab.II. We evaluated the changes in AP across different detection models, and the detection results are presented in Tab.III.
TABLE II: AEPatch under different settings of ${{\text{w}}_{0}}$ and ${{\text{w}}_{1}}$.
@@
TABLE III: The results of the comparative experiments between the two sets.

### v2

@@ IV-G Ablation Studies
To validate the effectiveness of the proposed dynamically optimized ensemble model, we conducted the following two experiments compared to the Average Ensemble model. We designed ensemble models utilizing the one-stage algorithms YOLOv2 and YOLOv3 and obtained respective adversarial patches, DOEPatch(YY) and AEPatch(YY). Additionally, we conducted experiments targeting Faster R-CNN and YOLOv3, which yielded novel adversarial patches, AEPatch(YF) and DOEPatch(YF). The weight parameters, denoted by ${{\text{w}}_{0}}$ and ${{\text{w}}_{1}}$, represent the weights before the detection model, which are constantly optimized during training in the dynamically optimized ensemble model. We manually adjusted the weight parameters of the average ensemble model and conducted five sets of experiments on AEPatch(YF), marked with numerical identifiers, as shown in Tab.VI. We evaluated the changes in AP across different detection models, and the detection results are presented in Tab.VII.
TABLE VI: AEPatch under different settings of ${{\text{w}}_{0}}$ and ${{\text{w}}_{1}}$.
@@
TABLE VII: The results of the comparative experiments between the two sets.
@@ V Conclusion
Acknowledgment
This work is funded by the National Natural Science Foundation of China (No.62103330, 62233014).